The area of a rectangle is 54 square units. Find the length and the width of the rectangle.
L = x+2 W = x-1
My work so far leads to me find that this unable to be factored to solve for x
(x+2)(x-1)=54
x2+2x-x-2=54
x2+x-2=54
x2+x-2-54=0
x2+x-56=0
(x-8)(x+7)=0
x=8 x=-7
can't use -7 so omit
x=8
back to original problem/formula
x+2 would read 8+2 = 10
x-1 would read 8-1= 7
IS THIS CORRECT??
Thank you for your help!

Answer by checkley77(12844)

(x+2)(x-1)=54
x2+2x-x-2=54
x2+x-2=54
x2+x-2-54=0
x2+x-56=0
(x-8)(x+7)=0 here's where you went wrong.
-----------------------------------------
(x-7)(x+8)=0
x-7=0
x=7
L=7+2=9
W=7-1=6
Proof:
9*6=54
---------------------------
x=8 x=-7
can't use -7 so omit
x=8